# A Systematic Review of Aspect-based Sentiment Analysis (ABSA): Domains, Methods, and Trends thanks: Preprint manuscript

### v4

@@ Abstract
Aspect-based Sentiment Analysis (ABSA) is a fine-grained type of sentiment analysis that identifies aspects and their associated opinions from a given text. With the surge of digital opinionated text data, ABSA gained increasing popularity for its ability to mine more detailed and targeted insights. Many review papers on ABSA subtasks and solution methodologies exist, however, few focus on trends over time or systemic issues relating to research application domains, datasets, and solution approaches. To fill the gap, this paper presents a Systematic Literature Review (SLR) of ABSA studies with a focus on trends and high-level relationships among these fundamental components. This review is one of the largest SLRs on ABSA, and also, to our knowledge, the first that systematically examines the trends and inter-relations among ABSA research and data distribution across domains and solution paradigms and approaches. Our sample includes 519 primary studies screened from 4191 search results without time constraints via an innovative automatic filtering process. Our quantitative analysis not only identifies trends in nearly two decades of ABSA research development but also unveils a systemic lack of dataset and domain diversity as well as domain mismatch that may hinder the development of future ABSA research. We discuss these findings and their implications and propose suggestions for future research.
@@ 1 Introduction
In the digital era, a vast amount of online opinionated text is generated daily through which people express views and feelings (i.e. sentiment) towards certain subjects, such as user reviews, social media posts, and open-ended survey question responses [1]. Understanding the sentiment of these opinionated text data is essential for gaining insights into people’s preferences and behaviours and supporting decision-making across a wide variety of domains [2, 3, 4, 5, 6]. The analyses of opinionated text usually aim at answering questions such as “What subjects were mentioned?”, “What did people think of (a specific subject)?”, and “How are the subjects and/or opinions distributed across the sample?” (e.g. [7, 8, 9, 10]). These objectives, along with today’s enormous volume of digital opinionated text, require an automated solution for identifying, extracting and classifying the subjects and their associated opinions from the raw text.
Aspect-based sentiment analysis (ABSA) is one such solution. ABSA is a sub-field of Sentiment Analysis (SA), which is a core task of natural language processing (NLP). SA, also known as “opinion mining” [5, 6, 11, 12, 13], solves the problem of identifying and classifying given text corpora’s affect or sentiment orientation [14, 4] into polarity categories (e.g. “positive, neutral, negative”) [15, 16, 17], intensity/strength scores (e.g. from 1 to 5) [18], or other categories. The “identifying the subjects of opinions” part of the quest relates to the granularity of SA. Traditional SA mostly focuses on document- or sentence-level sentiment and thus assumes a single subject of opinions [19, 20]. In recent decades, the explosion of online opinion text has attracted increasing interest in distilling more targeted insights on specific entities or their aspects within each sentence through finer-grained SA [19, 20, 14, 21, 22]. This is the problem Aspect-based sentiment analysis (ABSA) aims to solve.
A complete ABSA solution that identifies aspects and their corresponding sentiments typically consists of a combination of subtasks detailed in Subsection 2.1.1. Subtask combinations can, to a large extent, reflect different ABSA problem formulations and tactics. The techniques of solving these subtasks evolved from linguistic and statistical to the dominant machine learning (ML) approaches [23, 24, 20, 25], with Deep Learning (DL) methods on the rise in recent years [20, 26].
1.1 Review Rationale
This review is underpinned by the following rationales:
First, the shift towards ML, especially supervised and/or DL solutions for ABSA, highlights the importance of dataset resources, particularly annotated, large benchmark datasets for the quality and development of ABSA research. Meanwhile, the finer granularity of ABSA also brings the persistent challenge of domain dependency described in Subsection 2.1.2. The diversity of ABSA datasets and their domains can have a direct and systematic impact on research and applications.
The early seminal works in ABSA were motivated by commercial applications and focused on product and service reviews [20, 27, 26], such as Ganu et al., (2009) [28], Hu and Liu (2004a [29], 2004b [30]), and Pontiki et al. (2014 [31], 2015 [32], 2016 [33]) that laid influential foundations with widely-used product and service review ABSA benchmark datasets [27, 26]. However, the need for mining insights from opinions far exceeds this single domain. Many other areas, especially the public sector, also have an abundance of opinionated text data and can benefit from ABSA, such as helping policy-makers understand public attitudes and reactions towards events or changes [2], improving healthcare services and treatments via patient experience and concerns in clinical visits, symptoms, drug efficacy and side-effects [34, 35], and guiding educators in meeting teacher and learner needs and improving their experience [3, 4, 5, 6]. While the more general SA research has been applied to “nearly every domain” ( [19], P.1), this does not seem to be the case for ABSA. Chebolu et al. (2022) reviewed 62 public ABSA datasets released between 2004 and 2020 covering “over 25 domains” ( [36], p.1). However, 53 out of these 62 datasets were reviews of restaurants, hotels, and digital products; only five were not related to commercial products or services, and merely one was on the public sector domain (university reviews). Will this dataset domain homogeneity be found with a larger sample of primary studies? Does this domain skewness reflect the concentration of ABSA research focus or merely the lack of dataset diversity? Answers to these questions could inform and shape future ABSA research through individual research decisions and community resource collaboration.
Second, there are many good survey papers on ABSA, most focused on introducing the methodological details of common ABSA subtasks and solutions (e.g. [23, 37, 27, 38, 39, 40] or specific approaches such as DL methods for ABSA (e.g. [20, 26, 41, 42]). However, many of these reviews only explored each subtask and/or technique individually, and few examined their combinations or changes over time. For example, although the above-listed reviews [20, 26, 41, 42] reported the rise of DL approaches in ABSA similar to that of NLP as a whole, it is unclear whether ABSA research was also increasingly dominated by the attention mechanism from the Transformer architecture [43] and pre-trained large language models since 2018 [44], and if linguistic and traditional ML approaches were still active. In addition, most of these surveys used a smaller and selected sample that could not support conclusions on trends. As the field matures, we believe it is necessary and important to examine trends and matters outside the problem solution itself, so as to inform research decisions, identify issues, and call for necessary community awareness and actions.
@@ 1.2 Review Focus and Research Questions
  1. RQ1. To what extent is ABSA research and its dataset resources dominated by the commercial (especially the product and service review) domain? What proportion of ABSA research focuses on other domains and dataset resources?
@@
  2. RQ2. What are the most common ABSA problem formulations via subtask combinations, and what proportion of ABSA studies only focus on a specific sub-task?
  3. RQ3. What is the trend in the ABSA solution approaches over time? Are linguistic and traditional machine-learning approaches still in use?
@@
This review makes a number of unique contributions to the ABSA research field: 1) It is one of the largest scoped SLRs on ABSA, with 519 eligible studies filtered from 4191 search results without time constraints; 2) To our knowledge, it is the first SLR that systematically examines the ABSA data resource distribution in relation to research application domains and methodologies; and 3) Our review methodology adopted an innovative automatic filtering process based on PDF-mining, which enhanced screening quality and reliability. Our quantitative results not only revealed trends in nearly two decades of ABSA research literature but also highlighted potential systemic issues that could limit the development of future ABSA research.
1.3 Organisation of This Review
In the following text, we introduce ABSA including its subtasks and main challenges in Section 2, describe our SLR procedures in Section 3 (“Methods”), and answer the research questions using the SLR results in Section 4 (“Results”). Finally, we discuss the key findings and acknowledge limitations in Sections 5 and 6 (“Discussion” and “Conclusion”). Appendix A contains full details of the Methods section, and Appendix B provides additional figures from the Results section.
2 Aspect-Based Sentiment Analysis (ABSA)
The following examples illustrate the ABSA terminologies:
As demonstrated above, the fine granularity makes ABSA more targetable and informative than document- or sentence-level SA. Thus, ABSA can precede downstream applications such as attribute weighting in overall review ratings (e.g. [58]), aspect-based opinion summarization (e.g. [59, 60, 61], and automated personalised recommendation systems (e.g. [62, 63]).
2.1 ABSA Subtasks and Challenges
2.1.1 ABSA Subtasks
{quoting}
[vskip=12pt]
With the rise of multi-task learning and deep learning [81], an increasing number of studies explore ABSA under an End-to-End (E2E) framework that performs multiple fundamental ABSA subtasks in one model, and some combine them into a single composite task [64, 79, 82]. These composite tasks are most commonly formulated as a sequence- or span-based tagging problem [64, 79, 80]. The most common composite tasks are: Aspect-Opinion Pair Extraction (AOPE), which directly outputs {aspect, opinion} pairs from text input [80, 83, 84] such as “⟨menu, great⟩” from Example 1; Aspect-Polarity Co-Extraction (APCE) [64, 85], which outputs {aspect, sentiment polarity} pairs such as “⟨menu, positive⟩”; Aspect-Sentiment Triplet Extraction (ASTE) [64, 79, 86, 66], which outputs {aspect, opinion, sentiment category} triplets, such as “⟨menu, great, positive⟩”; and Aspect-Sentiment Quadraplet Extraction (ASQE) [87] that outputs {aspect, opinion, aspect category, sentiment category} quadruplets, such as “⟨menu, great, general, positive⟩”.
As this review focuses on trends instead of detailed solutions and methodologies, we refer interested readers to existing review papers that provide a good summary of common ABSA subtask solutions and approaches, such as AE (e.g. [23, 27, 38, 39]), deep learning (DL) approaches for ASC (e.g. [40]), and DL methods for ABSA (e.g. [20, 26, 41, 42]).
2.1.2 Context and Domain Dependencies
The nature and the interconnection of these subtasks determine that ABSA is heavily domain- and context-dependent [19, 36, 88]. Domain refers to the ABSA task (training or application) topic domains, and context can be either the “global” context of the document or the “local” context from the text surrounding a target word token or word chunks. At least in English, the same word or phrase could mean different things or bear different sentiments depending on the context and often also the topic domains. For example, “a big fan” could be an electric appliance or a person, depending on the sentence and the domain; “cold” could be positive for ice cream but negative for customer service; and “DPS (damage per second)” could be either a gaming aspect or non-aspect in other domains. Thus, the ability to incorporate relevant context is essential for ABSA solutions; and those with zero or very small context windows, such as n-gram and Markov models, are rare in ABSA literature and can only tackle a limited range of subtasks (e.g. [89]). Moreover, although many language models (e.g. Bidirectional Encoder Representations from Transformers (BERT, [90]), Generative pre-trained transformers (GPT, [91]), Recurrent Neural Network (RNN)-based models) already incorporated local context from the input-sequence and/or general context through pre-trained embeddings, they still performed unsatisfactorily on some ABSA domains and subtasks, especially IAE, AE with multi-word aspects, AE and ACD on mixed-domain corpora, and context-dependent ASC [92, 21, 11, 88]. Many studies showed that ABSA task performance benefits from expanding the feature space beyond generic and input textual context, with domain-specific dataset/representations, and additional input features such as Part-of-Speech (POS) tags, syntactic dependency relations, lexical databases, and domain knowledge graphs or ontologies [88, 21, 11]. Nonetheless, annotated datasets and domain-specific resources are costly to produce and limited in availability, and domain adaptation, as one solution to this, has been an ongoing challenge for ABSA [93, 82, 19, 88].
The above highlights the critical role of domain-specific datasets and resources in ABSA solution quality, especially for supervised approaches. On the other hand, it suggests the possibility that the prevalence of dataset-reliant solutions in the field and a skewed ABSA dataset domain distribution could systemically hinder ABSA solution performance and generalisability [93, 94], and thus confining ABSA research and solutions close to the resource-rich domains and languages. This idea underpins this literature review’s motivation and research questions.
@@ 3 Methods
Following the guidance of Kitchenham and Charters (2007) [45], we conducted this SLR with pre-planned scope, criteria, and procedures highlighted below. The complete SLR methods and process are detailed in Appendix A.
@@ 3.1 Main Procedures
The reviewed primary studies were from four major peer-reviewed digital databases: ACM Digital Library, IEEE Xplore, Science Direct, and Springer Link. First, we manually searched and extracted 4191 database search results without publication-year constraints. Appendix Section A.1 provides more details of the search strategies and results. Next, we applied the inclusion and exclusion criteria listed in Table 1 via automatic and manual screening steps and identified 519 in-scope peer-reviewed research publications for the review. The complete screening process is summarised in Table A.2 in Section A.2. We then manually reviewed the in-scope primary studies and recorded data following a planned scheme. Lastly, we checked, cleaned, and processed the extracted data and performed quantitative analysis against our RQs.
@@
3.2 Sample Summary
Figure 1 shows the number of total reviewed vs. included studies across all publication years for the 4191 search results. The search results include studies published between 1995 and 2023 (N=1), although all of the pre-2008 ones (2 from the 90s, 8 from 2003-2006, 17 from 2007) were not ABSA-focused and were excluded during automatic screening. The earliest in-scope ABSA study in the sample was published in 2008, followed by a very sparse period until 2013. The numbers of extracted and in-scope publications have both grown noticeably since 2014, a likely result of the emergence of deep learning approaches, especially sequence models such as RNNs [44, 95]. We also provide a breakdown of the included studies by publication year and type in Figure B.1 in Appendix B.
Figure 1: Number of Studies by Publication Year: Total reviewed vs. Included (N=4191)

### v5

# A Systematic Review of Aspect-based Sentiment Analysis: Domains, Methods, and Trends

@@ Abstract
Aspect-based Sentiment Analysis (ABSA) is a fine-grained type of sentiment analysis that identifies aspects and their associated opinions from a given text. With the surge of digital opinionated text data, ABSA gained increasing popularity for its ability to mine more detailed and targeted insights. Many review papers on ABSA subtasks and solution methodologies exist, however, few focus on trends over time or systemic issues relating to research application domains, datasets, and solution approaches. To fill the gap, this paper presents a Systematic Literature Review (SLR) of ABSA studies with a focus on trends and high-level relationships among these fundamental components. This review is one of the largest SLRs on ABSA. To our knowledge, it is also the first to systematically examine the interrelations among ABSA research and data distribution across domains, as well as trends in solution paradigms and approaches. Our sample includes 727 primary studies screened from 8550 search results without time constraints via an innovative automatic filtering process. Our quantitative analysis not only identifies trends in nearly two decades of ABSA research development but also unveils a systemic lack of dataset and domain diversity as well as domain mismatch that may hinder the development of future ABSA research. We discuss these findings and their implications and propose suggestions for future research.
@@ 1 Introduction
In the digital era, a vast amount of online opinionated text is generated daily through which people express views and feelings (i.e. sentiment) towards certain subjects, such as user reviews, social media posts, and open-ended survey question responses [1]. Understanding the sentiment of these opinionated text data is essential for gaining insights into people’s preferences and behaviours and supporting decision-making across a wide variety of domains [2, 3, 4, 5, 6]. The analyses of opinionated text usually aim at answering questions such as “What subjects were mentioned?”, “What did people think of (a specific subject)?”, and “How are the subjects and/or opinions distributed across the sample?” (e.g. [7, 8, 9, 10]). These objectives, along with today’s enormous volume of digital opinionated text, require an automated solution for identifying, extracting and classifying the subjects and their associated opinions from the raw text. Aspect-based Sentiment Analysis (ABSA) is one such solution.
@@ 1.1 Review Focus and Research Questions
  • RQ2. What are the most common ABSA problem formulations via subtask combinations, and what proportion of ABSA studies only focus on a specific subtask?
This review makes a number of unique contributions to the ABSA research field: 1) It is one of the largest scoped SLRs on ABSA, with a main review and a Phase-2 targeted review of a combined 727 primary studies published in 2008-2024, selected from 8550 search results without time constraint. 2) To our knowledge, it is the first SLR that systematically examines the ABSA data resource distribution in relation to research application domains and methodologies; and 3) Our review methodology adopted an innovative automatic filtering process based on PDF-mining, which enhanced screening quality and reliability. Our quantitative results not only revealed trends in nearly two decades of ABSA research literature but also highlighted potential systemic issues that could limit the development of future ABSA research.
1.2 Organisation of This Review
In Section 2 (“Background”), we introduce ABSA and highlight the motivation and uniqueness of this review. Section 3 (“Methods”) outlines our SLR procedures, and Section 4 (“Results”) answers the research questions with the SLR results. We then discuss the key findings and acknowledge limitations in Sections 5 and 6 (“Discussion” and “Conclusion”).
For those interested in more details, Appendix A provides an in-depth introduction to ABSA and its subtasks. Appendix B describes the full details of our Methods, and additional figures from the Results are provided in AppendixC.
2 Background
2.1 ABSA: a fine-grained sentiment analysis
Aspect-based Sentiment Analysis (ABSA) is a sub-field of Sentiment Analysis (SA), which is a core task of natural language processing (NLP). SA, also known as “opinion mining” [5, 6, 11, 12, 13], solves the problem of identifying and classifying given text corpora’s affect or sentiment orientation [14, 4] into polarity categories (e.g. “positive, neutral, negative”) [15, 16, 17], intensity/strength scores (e.g. from 1 to 5) [18], or other categories. The “identifying the subjects of opinions” part of the quest relates to the granularity of SA. Traditional SA mostly focuses on document- or sentence-level sentiment and thus assumes a single subject of opinions [19, 20]. In recent decades, the explosion of online opinion text has attracted increasing interest in distilling more targeted insights on specific entities or their aspects within each sentence through finer-grained SA [19, 20, 14, 21, 22]. This is the problem ABSA aims to solve.
2.2 ABSA and its subtasks
ABSA involves identifying the sentiments toward specific entities or their attributes, called aspects. These aspects can be explicitly mentioned in the text or implied from the context (“implicit aspects”), and can be grouped into aspect categories [23, 14, 24, 25, 26, 27]. Appendix A.1 presents a more detailed definition of ABSA, including its key components and examples.
A complete ABSA solution as described above traditionally involves a combination of subtasks, with the fundamental ones [28, 29, 30, 31, 32] being Aspect (term) Extraction (AE), Opinion (term) Extraction (OE), and Aspect-Sentiment Classification (ASC), or in an aggregated form via Aspect-Category Detection (ACD) and Aspect Category Sentiment Analysis (ACSA).
The choice of subtasks in an ABSA solution reflects both the problem formulation and, to a large extent, the technologies and resources available at the time. The solutions to these fundamental ABSA subtasks evolved from pure linguistic and statistical solutions to the dominant machine learning (ML) approaches [24, 33, 20, 34], usually with multiple subtask models or modules orchestrated in a pipeline [35, 36]. More recently, the rise of multi-task learning brought an increase in End-to-end (E2E) ABSA solutions that can better capture the inter-task relations via shared learning [37], and many only involve a single model that provides the full ABSA solution via one composite task [29, 35, 38]. The most typical composite ABSA tasks include Aspect-Opinion Pair Extraction (AOPE) [36, 39, 40], Aspect-Polarity Co-Extraction (APCE) [29, 41], Aspect-Sentiment Triplet Extraction (ASTE) [29, 35, 42, 31], and Aspect-Sentiment Quadruplet Extraction/Prediction (ASQE/ASQP) [43, 44, 45, 46]. We provide a more detailed introduction to ABSA subtasks in Appendix A.2.
2.3 The context- and domain-dependency challenges
The nature and the interconnection of its components and subtasks determine that ABSA is heavily domain- and context-dependent [19, 47, 48]. Domain refers to the ABSA task (training or application) topic domains, and context can be either the “global” context of the document or the “local” context from the text surrounding a target word token or word chunks. At least in English, the same word or phrase could mean different things or bear different sentiments depending on the context and topic domains. For example, “a big fan” could be an electric appliance or a person, depending on the sentence and the domain; “cold” could be positive for ice cream but negative for customer service; and “DPS” (damage per second) could be either a gaming aspect or non-aspect in other domains. Thus, the ability to incorporate relevant context is essential for ABSA solutions; and those with zero or very small context windows, such as n-gram and Markov models, are rare in ABSA literature and can only tackle a limited range of subtasks (e.g. [49]).
Moreover, although many language models (e.g. Bidirectional Encoder Representations from Transformers (BERT, [50]), Generative pre-trained transformers (GPT, [51]), Recurrent Neural Network (RNN)-based models) already incorporated local context from the input-sequence and/or general context through pre-trained embeddings, they still performed unsatisfactorily on some ABSA domains and subtasks, especially Implicit AE (IAE), AE with multi-word aspects, AE and ACD on mixed-domain corpora, and context-dependent ASC [52, 21, 11, 48]. Many studies showed that ABSA task performance benefits from expanding the feature space beyond the generic and input textual context. This includes incorporating domain-specific dataset/representations and additional input features such as Part-of-Speech (POS) tags, syntactic dependency relations, lexical databases, and domain knowledge graphs or ontologies [48, 21, 11]. Nonetheless, annotated datasets and domain-specific resources are costly to produce and limited in availability, and domain adaptation, as one solution to this, has been an ongoing challenge for ABSA [53, 38, 19, 48, 54].
The above highlights the critical role of domain-specific datasets and resources in ABSA solution quality, especially for supervised approaches. On the other hand, it suggests the possibility that the prevalence of dataset-reliant solutions in the field, and a skewed ABSA dataset domain distribution, could systemically hinder ABSA solution performance and generalisability [53, 55], thus confining ABSA research and solutions close to the resource-rich domains and languages. This idea underpins this literature review’s motivation and research questions.
2.4 Review Rationale
This review is motivated by the following rationales:
First, the shift towards ML, especially supervised and/or DL solutions for ABSA, highlights the importance of dataset resources. In particular, annotated large benchmark datasets are crucial for the quality and development of ABSA research. Meanwhile, the finer granularity of ABSA also brings the persistent challenge of domain dependency described in Subsection 2.3. The diversity of ABSA datasets and their domains can have a direct and systematic impact on research and applications.
The early seminal works in ABSA were motivated by commercial applications and focused on product and service reviews [20, 56, 57], such as Ganu et al., (2009) [58], Hu and Liu (2004a [59], 2004b [60]), and Pontiki et al. (2014 [61], 2015 [62], 2016 [63]) that laid influential foundations with widely-used product and service review ABSA benchmark datasets [56, 57]. Nevertheless, the need for mining insights from opinions far exceeds this single domain. Many other areas, especially the public sector, also have an abundance of opinionated text data and can benefit from ABSA, such as helping policy-makers understand public attitudes and reactions towards events or changes [2], improving healthcare services and treatments via patient experience and concerns in clinical visits, symptoms, drug efficacy and side-effects [64, 65], and guiding educators in meeting teacher and learner needs and improving their experience [3, 4, 5, 6]. While the more general SA research has been applied to “nearly every domain” [19, p. 1], this does not seem to be the case for ABSA. Chebolu et al. (2022) reviewed 62 public ABSA datasets released between 2004 and 2020 covering “over 25 domains” [47, p. 1]. However, 53 out of these 62 datasets were reviews of restaurants, hotels, and digital products; only five were not related to commercial products or services, and merely one was on the public sector domain (university reviews).
The above-mentioned evidence raises questions: Will this dataset domain homogeneity be found with a larger sample of primary studies? Does this domain skewness reflect the concentration of ABSA research focus or merely the lack of dataset diversity? This motivated our RQ1 (“To what extent is ABSA research and its dataset resources dominated by the commercial (especially the product and service review) domain? What proportion of ABSA research focuses on other domains and dataset resources?”) Answers to these questions could inform and shape future ABSA research through individual research decisions and community resource collaboration.
Second, there are many good survey papers on ABSA, most focused on introducing the methodological details of common ABSA subtasks and solutions (e.g. [24, 66, 56, 67, 68, 69] or specific approaches such as DL methods for ABSA (e.g. [20, 57, 70, 71, 72, 73, 54]). We list these surveys in Appendix A.3 as additional resources for the reader. Nonetheless, many of these reviews only explored each subtask and/or technique individually and often by iterating through reviewed studies, and few examined their combinations or changes over time and with quantitative evidence. For example, although the above-listed reviews [20, 57, 70, 71] reported the rise of DL approaches in ABSA similar to that of NLP as a whole, it is unclear whether ABSA research was also increasingly dominated by the attention mechanism from the Transformer architecture [74] and pre-trained large language models since 2018 [75], and if linguistic and traditional ML approaches were still active. In addition, most of these surveys used a smaller and selected sample that could not support conclusions on trends. As the field matures, we believe it is necessary and important to examine trends and matters outside the problem solution itself, so as to inform research decisions, identify issues, and call for necessary community awareness and actions. We thus proposed RQ2 (“What are the most common ABSA problem formulations via subtask combinations, and what proportion of ABSA studies only focus on a specific sub-task?”) and RQ3 (“What is the trend in the ABSA solution approaches over time? Are linguistic and traditional machine-learning approaches still in use?”).
@@ 2.4 Review Rationale
In summary, this review aims to address gaps in the ABSA literature. The high-level nature of our research questions is best answered through a large-scale SLR to provide solid evidence. The next section presents our SLR approach and sample.
@@ 3 Methods
Following the guidance of Kitchenham and Charters (2007) [76], we conducted this SLR with pre-planned scope, criteria, and procedures highlighted below. The complete SLR methods and process are detailed in Appendix B.
@@ 3.1 Main Procedures
For the main SLR sample, we sourced the primary studies in October 2022 from four major peer-reviewed digital databases: ACM Digital Library, IEEE Xplore, Science Direct, and SpringerLink. First, we manually searched and extracted 4191 database results without publication-year constraints. Appendix B.1 provides more details of the search strategies and results. Next, we applied the inclusion and exclusion criteria listed in Table 1 via automatic [^1] Our PDF mining for automatic review screening code is available at https://doi.org/10.5281/zenodo.12872948 and manual screening steps and identified 519 in-scope peer-reviewed research publications for the review. The complete screening process, including that of the automatic screening, is described in Appendix B.2. We then manually reviewed the in-scope primary studies and recorded data following a planned scheme. Lastly, we checked, cleaned, and processed the extracted data and performed quantitative analysis against our RQs.
@@
3.2 Main SLR Sample Summary
Figure 1 shows the number of total reviewed vs. included studies across all publication years for the 4191 SLR search results. The search results include studies published between 1995 and 2023 (N=1), although all of the pre-2008 ones (2 from the 90s, 8 from 2003-2006, 17 from 2007) were not ABSA-focused and were excluded during automatic screening. The earliest in-scope ABSA study in the sample was published in 2008, followed by a very sparse period until 2013. The numbers of extracted and in-scope publications have both grown noticeably since 2014, a likely result of the emergence of deep learning approaches, especially sequence models such as RNNs [75, 83]. We also present a breakdown of the included studies by publication year and type in Figure C.1 in Appendix C.
Figure 1: Number of Studies by Publication Year: Total reviewed (N=4191) vs. Included (N=519)